given the linear equation y=5x-1, find the y coordinates of the points (-1), (2), and (1,). Please show all of your work. Plot these points and graph the linear equation.
Answer by mananth(16946)

y=5x-1


(-1), (2), and (1,).
y=5*(-1)-1
y= -6 (-1,-6)

y=5*2-1
y=9 (2,9)
y=5*1-1
y=4 (1,4)